# Network Modelling of Asynchronous Change-Points in Multivariate Time Series

### v1

@@ 2.4 Hyperpriors
Motivated by our analysis of the hidden state process in Section 2.3, we now specify priors over the processes parameters $(\boldsymbol{W}_{0},\boldsymbol{W},\boldsymbol{A},\boldsymbol{q}_{0},\boldsymbol{q})$. A standard prior used to model random graphs is the Erdős-Rényi (ER) model (Erdős and Rényi,, 1959), which specifies $A_{i,j}\overset{\mathrm{iid}}{\sim}Bern(\rho)$ for $\rho\in(0,1)$. This prior assigns equal probability, $\rho$, to each edge being present in the graph. Our analysis in Section 2.3 suggests that the presence of cycles in the graph may result in undesirable behaviour in the change-point locations and so we would like our prior to assign little or no mass on such graphs.
Constructing a prior which assigns non-zero mass only to directed acyclic graphs is difficult due to the space being combinatorially complex and highly constrained by acyclicity. We instead base our prior on a modification of the ER model which does not allow self edges or parallel edges. Specifically, we let $\rho\in(0,1)$, and define our prior on $\boldsymbol{A}$ is as
@@
$Pr(A_{i,j}=a_{i,j},A_{j,i}=a_{j,i}|\rho)=\begin{cases}\rho/2,\;\;\;\;\;\;\;\text{ if }a_{i,j}=0,a_{j,i}=1\\ \rho/2,\;\;\;\;\;\;\;\text{ if }a_{i,j}=1,a_{j,i}=0\\ 1-\rho,\;\;\;\;\;\text{ if }a_{i,j}=0,a_{j,i}=0\\ 0,\;\;\;\;\;\;\;\;\;\;\;\text{ if }a_{i,j}=1,a_{j,i}=1\\ \end{cases}$ (21)
In other words, for each pair of distinct vertices $(i,j)$ with $i\neq j$, a directed edge is included with probability $\rho$, and omitted with probability $1-\rho$. Under this choice, the expected number of edges is $\rho d(d-1)/2$, allowing us to control the sparsity of the graph through our prior on $\rho$. To encourage sparse graphs, we place a uniform prior with parameters $(0,0.2)$ on $\rho$.
@@
where $Ga(a,b)$ denotes the Gamma distribution with shape $a$ and rate $b$. We find that this choice works well when modelling the dependence between pairs of lead-lag series. Under this choice, we are centering the weights around one, producing correlations in the lagged change-point locations less than 0.5, see Figure 1(c). This choice has the added property that the induced prior on the vector of normalised weights for series $j$ is
@@ 3 Posterior Inference
We now describe how to perform posterior inference for the NetCP model introduced in Section 2. Our MCMC sampler extends the work of Fearnhead, (2006) and Whiteley et al., (2011) to the multivariate setting. We show how the hidden states for each series can be efficiently sampled in blocks using the forward-backward recursions of Fearnhead, (2006). For long time series, we demonstrate how particle MCMC methods (Andrieu et al.,, 2010; Whiteley et al.,, 2011) can be used to control computational cost while preserving a high level of sampling efficiency.
The full joint posterior density is given by
$\pi(\boldsymbol{x},\boldsymbol{W}_{0},\boldsymbol{W},\boldsymbol{A},\boldsymbol{q}_{0},\boldsymbol{q},\rho|\boldsymbol{y})\propto f(\boldsymbol{y}|\boldsymbol{x})Pr(\boldsymbol{X}=\boldsymbol{x}|\boldsymbol{W}_{0},\boldsymbol{W},\boldsymbol{A},\boldsymbol{q}_{0},\boldsymbol{q})f(\boldsymbol{W}_{0})f(\boldsymbol{W})f(\boldsymbol{A}|\rho)f(\rho)f(\boldsymbol{q}_{0})f(\boldsymbol{q})$ (25)
The main difficulty arises when sampling the hidden states, $\boldsymbol{X}$. One approach is to sample each $X_{j,t}$ one at a time from its full posterior conditional (see Barry and Hartigan,, 1993; Quinlan et al.,, 2024). Since each $X_{j,t}$ is either $t-1$ (a change-point) or $X_{j,t-1}$ (no change-point), this can be performed with simple Bernoulli samples. While easy to implement, strong dependencies between the $X_{j,t}$ mean that this method tends to mix poorly (see Fan and Mackey,, 2017). Another popular approach is to use reversible jump MCMC (Green,, 1995), but again this is sub-optimal since designing efficient proposal moves is non trivial.
We construct an efficient sampling scheme by sampling the hidden states in each series as a block. This may be done using the method of Fearnhead, (2006), which employs a set of forward-backward recursions. However, these recursions have a computational cost of $\mathcal{O}(T^{2})$ per series, making them impractical within an MCMC framework. To balance sampling efficiency with computational cost, we instead use the particle approximation of Whiteley et al., (2011), maintaining the correct stationary distribution via conditional particle filtering within a particle Gibbs framework (Andrieu et al.,, 2010). A single iteration of our sampling scheme involves the following steps:
@@
    For $j=1,...,d$, sample $\boldsymbol{x}_{j,0:T}$ from $\pi(\boldsymbol{x}_{j,0:T}|\boldsymbol{y},\boldsymbol{x}_{(-j),0:T},\cdots)$ via a particle Gibbs step.
@@
We sample the static parameters in step 2 using standard Gibbs/Metropolis steps, the details of which can be found in Appendix A.
For the remainder of this section we discuss our blocked sampling scheme for the hidden states $\boldsymbol{x}_{j,:0:T}$ for $j=1,...,d$. Step 1 requires us to sample from the joint PMF of the hidden states in a given series $j$ that is,
$\pi(\boldsymbol{x}_{j,0:T}|\boldsymbol{y},\boldsymbol{x}_{(-j),0:T},\cdots)$ $\propto$ $Pr(\boldsymbol{X}_{j,0:T}=\boldsymbol{x}_{j,0:T}|\boldsymbol{X}_{(-j),0:T}=\boldsymbol{x}_{(-j),0:T})f_{j}(\boldsymbol{y}_{j,0:T}|\boldsymbol{x}_{j,0:T})$ (26)
Observe that due to the Markov properties of $\boldsymbol{X}_{t}$, the term $Pr(\boldsymbol{X}_{j,0:T}=\boldsymbol{x}_{j,0:T}|\boldsymbol{X}_{(-j),0:T}=\boldsymbol{x}_{(-j),0:T})$ can be factorised as
@@
where $\boldsymbol{x}^{[j=s]}=(x_{1},...,x_{j-1},s,x_{j+1},...,x_{d})^{T}$, denotes the vector $\boldsymbol{x}$ with the value $s$ imputed at the $j^{th}$ index. The hidden state at time $t$ can take two values ($t-1$ or $x_{j,t-1}$), and so the normalizing constant of $P_{t}^{(j)}(x_{j,t}|x_{j,t-1},\boldsymbol{x}_{(-j),0:T})$ can be easily found. Therefore, conditional on the hidden states in all other series, the hidden states in series $j$ follow Markov a chain, with transitions given by (28). Recall that the likelihood for series $j$ can be factorised as
@@
where $f_{j}(y_{j,t}|x_{j,t},\boldsymbol{y}_{j,0:t-1})$ is given by (4). The full posterior conditional (26) can therefore be written as
@@
where $x_{j,1}=0$ as per Equation (10).
The manner in which Equation (30) factorises allows us to use the forward-backward recursions of Fearnhead, (2006) to perform direct sampling. However, due to the $\mathcal{O}(T^{2})$ computational time needed for the exact recursions, we instead use a the of the method of Whiteley et al., (2011) to replace the exact filtering distributions at time $t$ with an approximation. This approximation consists of $N$ weighted particles and reduces the computational cost to $\mathcal{O}(NT)$. A key component in the method of Whiteley et al., (2011) is the particle filter of Fearnhead and Liu, (2007). It differs from standard particle methods in two ways. First, since the state space is discrete, a deterministic search of the proposal space is used instead of a random proposal, a strategy which greatly reduces Monte Carlo error. Second, instead of using multinomial resampling, it employs Stratified Optimal Resampling (SOR), which is shown to be optimal over all existing unbiased resampling methods in terms of the maximum Kolmogorov–Smirnov distance. This approach ensures that no duplicate particles are stored at each iteration, since there is no advantage in storing multiple copies in a discrete space (see Fearnhead and Liu,, 2007).
Algorithm 1 describes the resulting particle Gibbs update for $\boldsymbol{x}_{j,0:T}$. In the following, we denote the set of particles at time $t$ by $\mathcal{S}_{t}\subseteq\{0,...,t-1\}$ and their associated weights by $\mathcal{W}_{t}=\{w_{t}(s_{t})\}_{s_{t}\in S_{t}}$ Thus, at time $t$, the particle $s\in\mathcal{S}_{t}$ has weight $w_{t}(s)$.
@@
1:$\boldsymbol{x}^{*}$ (current hidden state matrix), $N$ (number of particles).
2:At time $t=1$, set $\mathcal{S}_{1}=\{0\}$ and $\mathcal{W}_{1}=\{1\}$.
3:for $t=2,...,T$ do
4:     Set $\mathcal{S}_{t}=\mathcal{S}_{t-1}\cup\{t-1\}$.
5:     For each $s_{t-1}\in\mathcal{S}_{t-1}$ set $w_{t}(s_{t-1})\propto w_{t-1}(s_{t-1})P_{t}^{(j)}(s_{t-1}|s_{t-1},\boldsymbol{x}_{(-j),0:T}^{*})f_{j}(y_{j,t}|s_{t-1},\boldsymbol{y}_{j,0:t-1})$
6:     Set $w_{t}(t-1)=f_{j}(y_{j,t})\sum\limits_{s_{t-1}\in\mathcal{S}_{t-1}}w_{t-1}(s_{t-1})P_{t}^{(j)}(t-1|s_{t-1},\boldsymbol{x}_{(-j),0:T}^{*})$
7:     Normalise the weights $\mathcal{W}_{t}$ such that $\sum\limits_{s_{t}\in\mathcal{S}_{t}}w_{t}(s_{t})=1$.
8:     If $|\mathcal{S}_{t}|>N$, resample $(\mathcal{S}_{t},\mathcal{W}_{t})$ using the Conditional SOR algorithm (conditional on $x_{j,t}^{*}$).
9:end for
10:Given $\{(\mathcal{S}_{t},\mathcal{W}_{t})\}_{t=1,...,T}$ generate a new sample $\boldsymbol{x}_{j,0:T}$ using the backward sampling algorithm.
Algorithm 1 requires additional algorithms to perform the resampling in step 7 and backward sampling in step 9. These algorithms are the same as those given in Whiteley et al., (2011) and can be found in the Appendix A.
Note that if the number of particles $N$ is set to $T$, then we never perform resampling in step 7 and Algorithm 1 is equivalent to the exact filtering recursions of Fearnhead, (2006). However, when $N<T$, the sampler maintains the correct stationary distribution due to conditioning on the previous sample (Andrieu et al.,, 2010). In our simulated and real datasets we find that relatively few particles ($N\approx 200$) are required to obtain mixing similar to that of the direct sampling method where $N=T$. An analysis of the efficiency of our particle Gibbs scheme compared to the single site Gibbs sampler can be found in Appendix B.
A key assumption in Algorithm 1 is that the segment specific parameters may be marginalised out of the model, producing a Rao-Blackwellized filtering scheme. This results in highly efficient Monte Carlo estimates although restricts us to specifying conjugate segment likelihood models. In cases where non-conjugate models are used, the particle filter of Fearnhead and Liu, (2007) may be replaced with that of Chopin, (2006). Here, the segment specific parameters may be left in the model and included in the filtering steps, albeit at an increased computational cost and a reduction in Monte Carlo efficiency.
Finally, we point out that it is trivial to modify the above particle Gibbs step to sample from the posterior of multivariate Markov chains other than the change-point model presented here. This is possible as long as the transition PMF is in the form given in (11) and the likelihood factorises according to (2). The updating in steps 3-5 of Algorithm 1 can be replaced with the more general particle filter of Fearnhead and Clifford, (2003), allowing for Markov models which could re-enter previous states.
@@ 4 Illustrations
In this section we evaluate the performance of NetCP in modelling change-points and inferring the underlying graph structure using both simulated and real data examples.
@@ 4.1 Simulation Study
    NetCP’s ability to perform edge link analysis on the true graph generating the change-points.
@@
    NetCP’s fit, in terms of log model evidence, against several competing methods.
We consider five scenarios for the change-point generating process and two possible segment likelihood models, resulting in 10 distinct data generating scenarios. Our change-point scenarios are designed to test situations in which change-points in different series (a) exhibit lead-lag relationships, (b) occur simultaneously across all, or subsets of the series and (c) are independent. Below we provide a description of each scenario S1 - S5.
@@
    Simulate from the NetCP model with parameters $\boldsymbol{W}_{0}=\boldsymbol{1}_{4}$ and $\boldsymbol{q}_{0}=\frac{1}{40}\boldsymbol{1}_{4}$. We use an adjacency matrix corresponding to a chain graph with $A_{1,2}=A_{2,3}=A_{3,4}=1$ and all other $A_{i,j}=0$. The edge weights are set to $W_{1,2}=W_{2,3}=W_{3,4}=5$ and the edge decay parameters to $q_{1,2}=q_{2,3}=q_{3,4}=0.6$. The correlations induced by this choice can be seen in Figure 1(c).
@@
    Simulate from the NetCP model with the same parameters as S1 but instead setting $A_{2,3}=0$, resulting in two sub-graphs.
@@
Many existing change-point models focus on detecting changes in means (Truong et al.,, 2020) and our second likelihood scenario is designed for this purpose. Under this scenario, the observations in each series are independent Gaussian with fixed variance $\sigma_{j}^{2}$ and piecewise constant means which are sampled independently from a $\mathcal{N}(0,\gamma_{j}^{2})$ distribution. This likelihood model is conjugate and the corresponding marginal segment density can be found in Appendix C. For this scenario’s simulated data we set $\sigma_{j}^{2}=0.5$ and $\gamma_{j}^{2}=3$ for all series $j$. For each scenario and likelihood model, we generate 50 datasets, each with $d=4$ time series and $T=500$ time points.
To evaluate NetCP’s ability to perform edge link analysis in test (i) we need to assess the model’s output for S1, S2 and S5 (where S5 corresponds to NetCP with an empty graph). Our aim is to check NetCP’s ability to correctly predict which edges $(i,j)$ are in the graph. For each simulated dataset we calculate the Area Under the Receiver Operating Characteristic Curve (AUC) by thresholding the posterior edge inclusion probabilities $Pr(A_{i,j}=1|\boldsymbol{y})$. The AUC is then averaged over the 50 datasets for each change-point scenario and likelihood model.
@@
S1 S2 S5
Gaussian Mean 0.994 (0.919, 1.0) 0.998 (0.998, 1.0) 1.0 (1.0, 1.0)
AR Process 0.987 (0.87, 1.0) 0.989 (0.899, 1.0) 1.0 (1.0, 1.0)
Table 1: Average AUC and 5% quantiles for edge link analysis under the NetCP model for change-point scenarios S1, S2 and S5.
Table 1 displays the AUC for the link analysis test. From this we can see that on a sample of $T=500$ time points, the model is able to nearly perfectly recover the true $4\times 4$ adjacency matrix from scenarios S1, S2 and S5.
For test (ii) we compare NetCP’s network prior to three competing change-point priors. The first is the special case of NetCP obtained by setting $\boldsymbol{A}=\boldsymbol{0}$, which is equivalent to the Bernoulli prior of Yao, (1984) being applied to each series independently. We refer to this model as the “Yao” model. We also consider the two priors proposed by Quinlan et al., (2024). The first is the Non-Global Correlated Change-Point PPM, which we refer to as “NG-CCP”. This prior is constructed by replacing the transition probabilities in (12) with
@@
where $t_{d}$ denotes the $d$-dimensional t distribution. We also consider the “Global” prior of Quinlan et al., (2024), which we refer to as “G-CCP”, obtained by setting
@@
$(\text{logit}(p^{*}_{1}),...,\text{logit}(p^{*}_{d}))$ $\sim$ $t_{d}(\nu,\mu,\Sigma)$ (34)
These priors encourage simultaneous change-points across series by allowing the change-point probabilities for a fixed time to be correlated through choice of $(\nu,\mu,\Sigma)$. Quinlan et al., (2024) do not provide priors for these parameters, however, they do provide a method to compute sensible default values. Both the G-CCP and NG-CCP models can be estimated using the particle Gibbs scheme described in Section 3 and the latent transition probabilities $p^{*}_{j,t}$ and $p^{*}_{j}$ sampled using the Metropolis steps given in Quinlan et al., (2024).
@@
The posterior predictive likelihoods $f_{M}(\boldsymbol{y}_{t}|\boldsymbol{y}_{0:t-1})$ can be approximated by running the Gibbs sampler using the data up to time $t-1$; at each iteration, the hidden state $x_{t}$ is sampled, and the predictive likelihood is estimated by averaging the likelihood of observation $\boldsymbol{y}_{t}$ over these sampled hidden states. To fit each model, we produce 5000 posterior samples using 150 particles and 500 burn-in. To achieve this we use the King’s College London (KCL) CREATE computing cluster.

### v2

@@ 2.4 Hyperpriors
Motivated by our analysis of the NetCP process in Section 2.3, we now specify priors over the processes parameters $(\boldsymbol{W}_{0},\boldsymbol{W},\boldsymbol{A},\boldsymbol{q}_{0},\boldsymbol{q})$. A standard prior used to model random graphs is the Erdős-Rényi (ER) model [59], which specifies a Bernoulli prior for the adjacency matrix elements, $A_{i,j}\overset{\mathrm{iid}}{\sim}Bern(\rho)$ for $\rho\in(0,1)$. This prior assigns equal probability, $\rho$, to each edge being present in the graph. However, our analysis in Section 2.3 suggests that cyclic graphs can result in undesirable behavior in the change-point locations (bursts of change-points) and so we want our prior to assign little or no mass on such graphs.
Constructing a prior which assigns non-zero mass only to directed acyclic graphs is difficult due to the space being combinatorially complex and highly constrained by acyclicity. We instead base our prior on a modification of the ER model which does not allow self edges or parallel edges. Specifically, we let $\rho\in(0,1)$, and define our prior on $\boldsymbol{A}$ as
@@
[…]
In other words, for each pair of distinct vertices $(i,j)$ with $i\neq j$, a single directed edge is included with probability $\rho$, and excluded with probability $1-\rho$. Under this choice, the expected number of edges is $\rho d(d-1)/2$, allowing us to control the sparsity of the graph through our prior on $\rho$. To encourage sparse graphs and prevent overfitting, we place a uniform prior with parameters $(0,0.2)$ on $\rho$.
@@
where $Ga(a,b)$ denotes the Gamma distribution with shape $a$ and rate $b$. We find that this choice works well when modeling the dependence between pairs of lead-lag series. Under this choice, we are centering the weights around one, producing correlations in the lagged change-point locations less than 0.5, see Figure 2. This choice has the added property that the induced prior on the vector of normalized weights for series $j$ is
@@ 3 Posterior Inference
We now describe how to perform posterior inference for the NetCP model introduced in Section 2. Our MCMC sampler extends the work of [62] and [90] to the multivariate setting. We show how the hidden states for each series can be efficiently sampled in blocks using the forward-backward recursions of [62]. For long time series, we demonstrate how particle MCMC methods [47, 90] can be used to control computational cost while preserving a high level of sampling efficiency. The full joint posterior density is given by
$\pi(\boldsymbol{x},\boldsymbol{W}_{0},\boldsymbol{W},\boldsymbol{A},\boldsymbol{q}_{0},\boldsymbol{q},\rho|\boldsymbol{y})$ $\propto f(\boldsymbol{y}|\boldsymbol{x})Pr(\boldsymbol{X}=\boldsymbol{x}|\boldsymbol{W}_{0},\boldsymbol{W},\boldsymbol{A},\boldsymbol{q}_{0},\boldsymbol{q})$ (29)
$\times f(\boldsymbol{W}_{0})f(\boldsymbol{W})f(\boldsymbol{A}|\rho)f(\rho)f(\boldsymbol{q}_{0})f(\boldsymbol{q}).$
The main difficulty arises when sampling the hidden states, $\boldsymbol{X}$. One approach is to sample each $X_{j,t}$ one at a time from its full posterior conditional [[, see]]Barry1993, Quinlan2024. Since each $X_{j,t}$ is either $1$ (a change-point) or $X_{j,t-1}+1$ (no change-point), this can be performed with simple Bernoulli samples. While easy to implement, strong dependencies between the $X_{j,t}$ mean that this method tends to mix poorly [[, see]]Fan2017. Another popular approach is to use reversible jump MCMC [67], but again this is sub-optimal since designing efficient proposal moves is non trivial.
We construct an efficient sampling scheme by sampling the hidden states in each series as a block. This may be done using the method of [62], which employs a set of forward-backward recursions. However, these recursions have a computational cost of $\mathcal{O}(T^{2})$ per series, making them impractical within an MCMC framework. To balance sampling efficiency with computational cost, we instead use the particle approximation of [90], maintaining the correct stationary distribution via conditional particle filtering within a particle Gibbs framework [47]. A single iteration of our sampling scheme involves the following steps:
@@
    For j=1,,dj=1,...,d, sample $\boldsymbol{x}_{j,0:T}$ from $\pi(\boldsymbol{x}_{j,0:T}|\boldsymbol{y},\boldsymbol{x}_{(-j),0:T},\cdots)$ via a particle Gibbs step.
@@
We sample the static parameters in step 2 using standard Gibbs/Metropolis steps, the details of which can be found in Appendix B.
For the remainder of this section we discuss our blocked sampling scheme for the hidden states $\boldsymbol{x}_{j,:0:T}$ for j=1,,dj=1,...,d. Step 1 requires us to sample from the joint PMF of the hidden states in a given series $j$ that is,
$\pi(\boldsymbol{x}_{j,0:T}|\boldsymbol{y},\boldsymbol{x}_{(-j),0:T},\cdots)$ $\propto$ $Pr(\boldsymbol{X}_{j,0:T}=\boldsymbol{x}_{j,0:T}|\boldsymbol{X}_{(-j),0:T}=\boldsymbol{x}_{(-j),0:T})f_{j}(\boldsymbol{y}_{j,0:T}|\boldsymbol{x}_{j,0:T}).$ (30)
Observe that due to the Markov properties of $\boldsymbol{X}_{t}$, the term $Pr(\boldsymbol{X}_{j,0:T}=\boldsymbol{x}_{j,0:T}|\boldsymbol{X}_{(-j),0:T}=\boldsymbol{x}_{(-j),0:T})$ can be factorized as
@@
where $\boldsymbol{x}^{[j=s]}=(x_{1},...,x_{j-1},s,x_{j+1},...,x_{d})^{T}$, denotes the vector $\boldsymbol{x}$ with the value $s$ imputed at the $j^{th}$ index. The hidden state at time $t$ can take two values ($1$ or $x_{j,t-1}+1$), and so the normalizing constant of $P_{t}^{(j)}(x_{j,t}|x_{j,t-1},\boldsymbol{x}_{(-j),0:T})$ can be easily found. Therefore, conditional on the hidden states in all other series, the hidden states in series $j$ follow Markov a chain, with transitions given by Equation (32). Recall that the likelihood for series $j$ can be factorized as
@@
where $f_{j}(y_{j,t}|x_{j,t},\boldsymbol{y}_{j,0:t-1})$ is given by Equation (4). The full posterior conditional, Equation (30), can therefore be written as
@@
where $x_{j,1}=1$ as per Equation (10).
The manner in which Equation (34) factorizes allows us to use the forward-backward recursions of [62] to perform direct sampling. However, due to the $\mathcal{O}(T^{2})$ computational time needed for the exact recursions, we instead use the method of [90] to replace the exact filtering distributions at time $t$ with an approximation. This approximation consists of $N$ weighted particles and reduces the computational cost to $\mathcal{O}(NT)$. A key component in the method of [90] is the particle filter of [64]. It differs from standard particle methods in two ways. First, since the state space is discrete, a deterministic search of the proposal space is used instead of a random proposal, a strategy which greatly reduces Monte Carlo error. Second, instead of using multinomial resampling, it employs Stratified Optimal Resampling (SOR), which is shown to be optimal over all existing unbiased resampling methods in terms of the maximum Kolmogorov–Smirnov distance. This approach also ensures that no duplicate particles are stored at each iteration, since there is no advantage in storing multiple copies in a discrete space [[, see]]Fearnhead2007.
Algorithm 1 describes the resulting particle Gibbs update for $\boldsymbol{x}_{j,0:T}$. In the following, we denote the set of particles at time $t$ by $\mathcal{S}_{t}\subseteq\{1,...,t\}$ and their associated weights by $\mathcal{W}_{t}=\{w_{t}(s_{t})\}_{s_{t}\in S_{t}}$. Therefore, at time $t$, the particle $s\in\mathcal{S}_{t}$ has weight $w_{t}(s)$.
@@
[…]
2:At time $t=1$, set $\mathcal{S}_{1}=\{1\}$ and $\mathcal{W}_{1}=\{1\}$.
3:for t=2,,Tt=2,...,T do
4:  Set $\mathcal{S}_{t}=\{s_{t-1}+1:s_{t-1}\in\mathcal{S}_{t-1}\}\cup\{1\}$.
5:  for each $s_{t-1}\in\mathcal{S}_{t-1}$ do
6:   Set $w_{t}(s_{t-1}+1)\propto w_{t-1}(s_{t-1})P_{t}^{(j)}(s_{t-1}+1|s_{t-1},\boldsymbol{x}_{(-j),0:T}^{*})f_{j}(y_{j,t}|s_{t-1}+1,\boldsymbol{y}_{j,0:t-1})$
7:  end for
8:  Set $w_{t}(1)=f_{j}(y_{j,t})\sum\limits_{s_{t-1}\in\mathcal{S}_{t-1}}w_{t-1}(s_{t-1})P_{t}^{(j)}(1|s_{t-1},\boldsymbol{x}_{(-j),0:T}^{*})$
9:  Normalize the weights $\mathcal{W}_{t}$ such that $\sum\limits_{s_{t}\in\mathcal{S}_{t}}w_{t}(s_{t})=1$.
10:  If $|\mathcal{S}_{t}|>N$, resample $(\mathcal{S}_{t},\mathcal{W}_{t})$ using the Conditional SOR algorithm (conditional on $x_{j,t}^{*}$).
11:end for
12:Given $\{(\mathcal{S}_{t},\mathcal{W}_{t})\}_{t=1,...,T}$ generate a new sample $\boldsymbol{x}_{j,0:T}$ using the backward sampling algorithm.
Algorithm 1 requires additional algorithms to perform the resampling in step 9 and backward sampling in step 11. These algorithms are the same as those given in [90] and can be found in Appendix B.
Note that if the number of particles $N$ is set to $T$, then we never perform resampling in step 9 and Algorithm 1 is equivalent to the exact filtering recursions of [62]. However, when $N<T$, the sampler maintains the correct stationary distribution due to conditioning on the previous sample [47]. In our simulated and real datasets we find that relatively few particles ($N\approx 200$) are required to obtain mixing similar to that of the direct sampling method where $N=T$. On a small simulated dataset, we find that the particle Gibbs with only $N=50$ obtains an integrated autocorrelation time that is 49 times smaller than that of the single-site Gibbs sampler which updates $X_{j,t}$ one at a time. Details of this analysis can be found in Appendix C.
A key assumption in Algorithm 1 is that the segment specific parameters may be marginalized out of the model, producing a Rao-Blackwellized filtering scheme. This results in highly efficient Monte Carlo estimates although restricts us to specifying conjugate segment likelihood models. The approach could be extended to handle NetCP models with non-conjugate segment densities by replacing the particle filter of [64] with that of [55]. Here, the segment specific parameters may be left in the model and included in the filtering steps, albeit at an increased computational cost and a reduction in Monte Carlo efficiency.
Finally, we point out that it is trivial to modify the above particle Gibbs step to sample from the posterior of multivariate Markov chains other than the change-point model presented here. This is possible as long as the transition PMF is in the form given in Equation (11) and the likelihood factorizes according to Equation (2). The updating in steps 3-7 of Algorithm 1 can be replaced with the more general particle filter of [63], allowing for Markov models which could re-enter previous states.
@@ 4 Illustrations
In this section we evaluate the performance of NetCP in modeling change-points and inferring the underlying graph structure using both simulated and real data examples.
@@ 4.1 Simulation Study
    NetCP’s ability to infer the true underlying graph structure.
@@
    NetCP’s fit, in terms of log model evidence, against several competing priors.
We consider five scenarios for the change-point generating process and two possible segment likelihood models, resulting in 10 distinct data generating scenarios. In all cases, we use dimension $d=4$. Our change-point scenarios are designed to test situations in which change-points in different series (a) exhibit lead-lag relationships, (b) occur simultaneously across all, or subsets of the series and (c) are independent. Below we provide a description of each scenario S1 - S5.
@@
    Simulate from the NetCP process with parameters $\boldsymbol{W}_{0}=\boldsymbol{1}_{4}$ and $\boldsymbol{q}_{0}=\frac{1}{40}\boldsymbol{1}_{4}$. We use an adjacency matrix corresponding to a chain graph with $A_{1,2}=A_{2,3}=A_{3,4}=1$ and all other $A_{i,j}=0$. The edge weights are set to $W_{1,2}=W_{2,3}=W_{3,4}=5$ and the edge decay parameters to $q_{1,2}=q_{2,3}=q_{3,4}=0.6$. The correlations induced by this choice can be seen in Figure 2.
@@
    Simulate from the NetCP process with the same parameters as S1 but instead setting $A_{2,3}=0$, resulting in two sub-graphs.
@@
Many existing change-point models focus on detecting changes in means [87] and our second likelihood scenario is designed for this purpose. Under this scenario, the observations in each series are independent Gaussian with fixed variance $\sigma_{j}^{2}$ and piecewise constant means which are sampled independently from a $\mathcal{N}(0,\gamma_{j}^{2})$ distribution. This likelihood model is conjugate and the corresponding marginal segment density can be found in Appendix D. For this scenario’s simulated data we set $\sigma_{j}^{2}=0.5$ and $\gamma_{j}^{2}=3$ for all series $j$. For each scenario and likelihood model, we generate 50 datasets, each with $d=4$ time series and $T=500$ time points.
To evaluate NetCP’s ability to infer the true underlying graph, we assess the model’s output for S1, S2, and S5 (S5 corresponds to NetCP with an empty graph). Specifically, we examine whether NetCP correctly predicts edges $(i,j)$ in the graph by comparing the posterior probabilities $Pr(A_{i,j}=1\mid\boldsymbol{y})$ to the true adjacency matrix for each scenario. Classification performance is quantified using the Area Under the Receiver Operating Characteristic (ROC) Curve (AUC), which plots the true positive rate against the false positive rate for varying thresholds of $Pr(A_{i,j}=1\mid\boldsymbol{y})$. AUC values are averaged over the 50 simulated datasets for each change-point scenario and likelihood model. However, for scenario S5, the true adjacency matrix contains only zeros (no edges), making the AUC undefined. In this case, we report the true negative rate with a threshold of 0.5, averaged over the 50 datasets.
@@
S1 S2
Gaussian Mean 0.99 (0.84, 1.00) 0.99 (1.00, 1.00)
AR Process 0.97 (0.71, 1.00) 0.97 (0.71, 1.00)
Table 1: Mean and 5% quantiles of the Area Under the ROC Curve (AUC) computed over 50 datasets for scenarios S1 and S2, under both AR process and Gaussian mean likelihoods. Higher AUC values indicate better edge prediction performance.
Table 1 displays the AUC for the edge prediction test on scenarios S1 and S2. For scenario S5, where AUC is undefined, we threshold the inclusion probabilities $Pr(A_{i,j}=1|\boldsymbol{y})$ at 0.5, obtaining an average true negative rate of 0.998 and 0.987 for the Gaussian mean and AR Process likelihoods respectively. From this we can see that on a sample of $T=500$ time points, NetCP is able to nearly perfectly recover the true $4\times 4$ adjacency matrix under each scenario.
For test (ii) we compare NetCP to three competing change-point priors. The first is the special case of NetCP obtained by setting $\boldsymbol{A}=\boldsymbol{0}$, which is equivalent to the Bernoulli prior of [92] being applied to each series independently. We refer to this model as the “Yao” model. We also consider the two priors proposed by [84]. The first is the Non-Global Correlated Change-Point PPM, which we refer to as “NG-CCP”. This prior is constructed by replacing the transition probabilities in Equation (12) with
@@
where $t_{d}$ denotes the $d$-dimensional t distribution. We also consider the “Global” prior of [84], which we refer to as “G-CCP”, obtained by setting
@@
$(\text{logit}(p^{*}_{1}),...,\text{logit}(p^{*}_{d}))$ $\sim$ $t_{d}(\nu,\mu,\Sigma).$ (38)
These priors encourage simultaneous change-points across series by allowing the change-point probabilities for a fixed time to be correlated through choice of $(\nu,\mu,\Sigma)$. [84] do not provide priors for these parameters, however, they do provide a method to compute sensible default values. Both the G-CCP and NG-CCP models can be estimated using the particle Gibbs scheme described in Section 3 and the latent transition probabilities $p^{*}_{j,t}$ and $p^{*}_{j}$ sampled using the Metropolis steps given in [84].
@@
The posterior predictive likelihoods $f_{M}(\boldsymbol{y}_{t}|\boldsymbol{y}_{0:t-1})$ can be approximated by running the Gibbs sampler using the data up to time t1t-1; at each iteration, the hidden state $x_{t}$ is sampled, and the predictive likelihood is estimated by averaging the likelihood of observation $\boldsymbol{y}_{t}$ over these sampled hidden states. To fit each model, we produce 5000 posterior samples using 150 particles and 500 burn-in. To achieve this we use the King’s College London (KCL) CREATE computing cluster.